OpenStudy (anonymous):

find the limit of :

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty} \sqrt x * (\sqrt(x+c) -\sqrt(x) )\]

OpenStudy (anonymous):

is this correct? \[\lim_{x \rightarrow \infty} \sqrt x*(\sqrt(x+c) -\sqrt(x) *\left( (\sqrt x+c)+ \sqrt(x) \right)\]

terenzreignz (terenzreignz):

You multiplied something to it, you better divide the whole expression by that which you multiplied it with so that it doesn't "hurt" your original expression.

OpenStudy (amistre64):

\[\lim_{x \rightarrow \infty} \sqrt x * (\sqrt(x+c) -\sqrt(x) )\] \[\lim_{x \to \infty} \frac{ \sqrt{x+c} -\sqrt x}{\frac{1}{x^{1/2}}}\]can you form it into a shape that can be Lhospitalized?